Liam Company signed a lease for an office building for a period of 12 years. Under the lease agreement, a security deposit of $8,200 is made. The deposit will be returned at the expiration of the lease with interest compounded at 5% per year.

What amount will Liam receive at the time the lease expires? (Round answer to 2 decimal places, e.g. 25.25.)
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

A=8200*(1.05)^12

=8200*1.79585633

=$14726.02(Approx).